traffic signs are regulated by the manual on uniform traffic control devices. The perimeter of some traffic signs must be 74 inches. Also, its length is 9 inches longer than its width.
Answer by VickyFM(4)

Perimeter= 2L+2W
We know the perimeter: "The perimeter of some traffic signs must be 74 inches."
74=2L+2W
The question also says: "Also, its length is 9 inches longer than its width." This implies that the length is longer so we use addition 9+W
We still don't know the width. There is nothing to plug into the 2W of the Perimeter equation.
The equation now looks like 74=2(9+W)+2W
74=18+2W+2W
74=18+4W
56=4W
W=14